OpenStudy (anonymous):

Evaluate: (-9) * 7 * (-18) =

OpenStudy (anonymous):

is that multiply?

OpenStudy (anonymous):

negative times a positive is negative; negative times a negative is positive: -9*7= -63 -63*(-18)= +1134

OpenStudy (anonymous):

ohhhh yup @ShaeWX beat me to that......... hehe

OpenStudy (anonymous):

Yeah, the asterisks don't show well

OpenStudy (anonymous):

idk what that is... but ok!

OpenStudy (anonymous):

yes multiply